for rice by 10%. Calculate her income elasticity of demand for rice using point elasticity metho


1
Expert's answer
2021-04-30T10:47:09-0400

Income elasticity of demand is calculated as:

"E_y=\dfrac{\Delta Q}{\Delta Y}\times \dfrac{Y}{Q}"

Or:

"E_y=\dfrac{\%\Delta Q}{\%\Delta Y}"